# PyraTrans: Learning Attention-Enriched Multi-Scale Pyramid Network from Pre-Trained Transformers for Effective Malicious URL Detection

### 01 Dec 2023 v1

@@ Abstract
Detecting malicious URLs is a crucial aspect of web search and mining, significantly impacting internet security. Though advancements in machine learning have improved the effectiveness of detection methods, these methods still face significant challenges in their capacity to generalize and their resilience against evolving threats. In this paper, we propose PyraTrans, an approach that combines the strengths of pretrained Transformers and pyramid feature learning for improving malicious URL detection. We implement PyraTrans by leveraging a pretrained CharBERT as the base and augmenting it with 3 connected feature modules: 1) The Encoder Feature Extraction module, which extracts representations from each encoder layer of CharBERT to obtain multi-order features; 2) The Multi-Scale Feature Learning Module, which captures multi-scale local contextual insights and aggregate information across different layer-levels; and 3) The Pyramid Spatial Attention Module, which learns hierarchical and spatial feature attentions, highlighting critical classification signals while reducing noise. The proposed approach addresses the limitations of the Transformer in local feature learning and spatial awareness, and enabling us to extract multi-order, multi-scale URL feature representations with enhanced attentional focus. PyraTrans is evaluated using 4 benchmark datasets, where it demonstrated significant advancements over prior baseline methods. Particularly, on the imbalanced dataset, our method, trained with 60% of the data, achieves a TPR increase of 16-24 points and an F1-score improvement of 12-27 points. With just 10% of the data for training, the TPR is 3.3-6.5 times and the F1-score is 2.9-4.5 times that of the baseline. Additionally, our approach not only demonstrates robustness against adversarial sample attacks, achieving an accuracy of 0.9023, significantly surpassing the best baseline accuracy of 0.7610, but it shows improved generalization in cross-dataset evaluation experiments. We also executed case studies on 30 active malicious websites, where our technique achieved 100% accuracy in identifying all malicious sites. Codes and data are available at: https://github.com/Alixyvtte/PyraTrans.
@@
Cybersecurity, Phishing Detection, Transformer, Pre-trained Language Model, Deep Learning
@@ I Introduction
Recently, pretrained Transformer frameworks, such as BERT [15], have demonstrated promising prospects in various fields including natural language understanding, computer vision, and bioinformatics. This novel computational architecture and training paradigm offer new opportunities in malicious URL detection [16, 17]. They not only have stronger contextual learning capabilities but are also purely data-driven, end-to-end models. Table I presents an example of a BERT token sequence generated from a URL. However, despite the success of pretrained Transformer across multiple disciplines, their technological advantages have yet to be fully explored and manifested in malicious URL detection.
@@
In this study, we propose PyraTrans, a novel appraoch developed on the foundation of pretrained Transformers, aimed at effectively detecting malicious URLs. This method exploits the various encoding layers of pre-trained Transformers and amalgamates them with multi-scale feature extraction and pyramid attention. Our approach, when tested across a broad range of public benchmarks datasets, exhibited superior performance over previous methods, proving to be more generalizable and robust, particularly in challenging situations like adversarial sample attacks, cross-dataset evaluation, class-imbalanced learning. We also conduct a case study to evaluate the model in real-world applications.
@@
  • Our method establishes new state-of-the-art performance in malicious URL detection, achieving over a 10% performance increase across various test scenarios compared to current state-of-the-art methods. In certain challenging scenarios, the improvement in the F1-score even reaches 40 points.
  • Our approach leverages the multi-layer encoding capability of Transformers to construct a multi-order URL representation. Through empirical analysis, we have found that this layered Transformer encoding significantly improves performance in detecting malicious URLs.
  • By embedding multi-scale feature learning and pyramid attention into the pre-trained Transformer architecture, our method rectifies the inherent limitations of Transformers in terms of local feature extraction and spatial relationship perception.
  • Our research exposes weaknesses in generalization and robustness in previous methods. Particularly, our case study demonstrates a significant decrease in the detection success rate of these methods when applied in practical applications. In contrast, our approach has improved performance in these critical aspects.
@@ IV Methodology
We propose a method that incorporates a backbone network constructed by the CharBERT model, along with three feature modules: the Encoder Feature Extraction module, the Multi-scale Learning module, and the Pyramid Spatial Attention module.
@@
  • Backbone Network: The CharBERT model, acting as the backbone network, improves URL data interpretation and analysis with its advanced subword and character-level embedding which extends based on BERT.
  • Encoder Feature Extraction: This module aggregates and extracts multi-level features, employing a novel feature integration approach to effectively capture and utilize both high-level and low-level features from encoders.
  • Multi-scale Learning: This module is designed to capture features of different scales and granularities and extract both local and global features from URL strings. While maintaining a lightweight structure, it adaptively recalibrates the information flow from multiple channels, enhancing the model’s focus on key data aspects.
  • Pyramid Spatial Attention: This module emphasizes the contextual relevance of various layers within the network. It enhances the base network by laterally incorporating Spatial Pyramid Attention (SPA) Blocks, prioritizing specific features over others based on their contextual importance.
@@
Figure 3: PyraTrans: Consists of four essential components. CharBERT, the backbone network for feature extraction from characters and subwords; Encoder feature extraction and concatenation from different layers; A multi-scale learning module is followed by a spatial attention mechanism that adaptively reweights various informative feature channels, enhancing the classifier’s ability to perform classification.
@@ IV-A Backbone Network
We employ the pretrained CharBERT[28] as our backbone network, chosen for its simultaneous focus on subword and character-level features of text. CharBERT, an extension of the BERT model, integrates the Transformer architecture with a novel dual-channel framework designed to capture information at both subword and character levels. CharBERT achieves notable progress through the implementation of two key components: (1) the Character Embedding Module, encoding character sequences from input tokens and (2) the Heterogeneous Interaction Module, combines features from both character and subword channels, and then independently separates them into distinct representations as input for the encoder layer.
Within the character embedding module, we generate token-level embeddings by incorporating character-level information into input sentences. The character-aware embedding of each token is primarily generated through two components: the encoding of individual characters and subword units. These two components are integrated via a dual-channel architecture. To establish contextual character embeddings, we utilize a bidirectional Gated Recurrent Unit (BiGRU) layer. The BiGRU employs a bidirectional recurrent neural network with only the input and forget gates[29]. The architecture diagram of the BiGRU is depicted in Figure 2.
@@ IV-B Feature Extraction
However, in the context of the BERT model, even though each layer of the model takes the previous layer’s output features as input for computation, multiple intricate calculations within each layer’s processing may still result in potential degradation of lower-level and mid-level features, which is detrimental to the complete feature learning process. Li et al. [32] use feature concatenation to integrate aspect features from each layer of BERT for aspect term sentiment classification. This approach, instead of relying only on the final layer for classification features, effectively enhances classification performance by leveraging the distinct features learned at each layer of BERT.
@@
Our method uses a hierarchical feature extraction model to exploit the multi-level information from the low, middle, and high layers of CharBERT. The feature extraction process is as follows:
Consider a sequence of outputs $k_{1},k_{2},...,k_{n}$ and $u_{1},u_{2},...,u_{m}$, where each output $k_{i}$ and $u_{j}$ has a rank of $(H,W,C)$, representing the outputs of CharBERT’s word-level and character-level encoders at various layers. $H$ is the batch size, $W$ is the fixed URL sequence length (200 in our model), and $C$ is a 768-dimensional vector for each merged hidden layer output in CharBERT. For example, $k_{1}$ and $u_{2}$ be two tensors representing the sequence and character embeddings, respectively. Let $w$ be the sequence length, and $d$ be the embedding dimension:
@@ IV-C Multi-scale Learning
After feature concatenation, we introduce a lightweight multi-scale module for processing text embedding information across various scales. Initially, dilated convolutions with distinct dilation rates are employed to capture multi-scale information, allowing for a larger receptive field without increasing parameters. Subsequently, we utilize depthwise separable convolutions (referred to as DSConv)[33] to enhance efficiency by reducing floating-point operations and model parameters. These convolutions serve as fundamental operators for extending network depth and width. Formally, let the input feature map be denoted as $M\in\mathbb{R}^{C\times H\times W}$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width. The process begins by applying a single DSConv (conv3 x 3) to $M$, extracting common information denoted as $F_{0}$ for each branch. Specifically:
@@
IV-D Pyramid Spatial Attention
In the multi-scale learning module, a simple element-wise summation of features from different scales may inadvertently downplay the importance of informative branches, treating all scales equally. To address this issue, we employ a spatial pyramid attention mechanism [35], which evaluates subfields across multiple scales, adjusts branch weights, and enhances overall performance.
@@ V Experiments
    Evaluation of configurations of stacked feature layers for multi-layer feature extraction.
@@
    Investigation of model data scale dependency through experiments with diverse training dataset sizes.
@@
    Varying the size of training datasets to assess the model’s sensitivity to data scale.
@@
    Evaluation of multi-class classification performance.
@@
    Robustness assessment through adversarial sample testing.
@@
    Examining instances of the lastest active malicious URLs to evaluate the practicality.
Setup. Considering the scale of the detection task and computational resources, we opted for the pre-trained CharBERT based on BERT. This model was trained on the English Wikipedia corpus, consisting of a total of 12GB and approximately 2,500 million words. Hyperparameter tuning during fine-tuning led to a batch size of 64, AdamW optimizer (initial learning rate: 2e-5, weight decay: 1e-4), 0.1 dropout rate, and 5 training epochs. We utilized PyTorch 2.0, NVIDIA CUDA 11.8, and Python 3.8, conducting training on NVIDIA A100 GPUs. The final model for each experiment was selected based on the best validation loss.
@@
V-A Multi-Layer Feature Evaluation
We developed a multi-layer feature extraction module to distill semantic features from different layers of the Transformer encoders in our backbone network. Our approach involves stacking embedding outputs from different layers to form a multi-layered feature matrix, which we optimized to enhance URL semantic feature representation. Through this, we investigated the complexity of feature representation across the network and aimed to demonstrate the effectiveness of multi-layered features in improving overall model performance.
We created a training corpus from the GramBeddings dataset, randomly selecting 128k URLs from the 80k available in the training set while maintaining the proportion of malicious and benign URLs. We also randomly sampled 32k URLs for testing and validation respectively. The performance of various stacked configurations and the incremental gains shown in Table IV. We observed a consistent improvement in evaluation metrics as we incrementally incorporated additional layers, namely the last 2, 3, 4, and 5 layers. Ultimately, the stacking of 12 embedding output layers achieved the best performance in URL detection tasks, demonstrating the effectiveness of integrating both lower and deeper layers in the model architecture.
V-B Comparative Study
In this section, we compare the performance of PyraTrans with several proposed baselines using binary and multi-class malicious URL detection datasets.
@@ V-B1 Binary classification
As shown in Table V, our method outperforms other methods across several metrics in a class-imbalanced scenario. Even with only 1% of training data, our model attains a 0.5 F1 score, which is four times higher than the baseline’s performance. With more training data, our F1 score reaches 0.7027, which is much higher than the baseline’s peak of 0.5515. Our model also demonstrates excellent precision and recall, and a lower false negative rate is crucial in the field of malicious URL detection, as we aim to minimize missing malicious URLs. The AUC of 0.9438 showcases the overall performance of our model on this unbalanced dataset.
We note that URLNet suffers from a high false positive rate, especially with smaller datasets, despite achieving high precision on larger datasets. Due to its reliance on the size and diversity of the training set, Grambeddings also encounters challenges. Our method shows consistent and reliable performance across diverse data distributions, indicating its ability to differentiate malicious and benign samples, and overcoming the challenges of imbalanced datasets better than other methods.
V-B2 Multiple classification
@@
Fig.6 illustrates the performance of our method and the baseline methods across all four categories. Our method surpasses the baseline methods in each metric. The average ROC curve highlights our method’s efficacy with a TPR of almost 90% at a low FPR of 0.001, surpassing other methods that achieve around 75%. Grambeddings struggled in recognizing negative samples from various categories, resulting in a 50% F1 score for defacement and phishing URLs and an overall accuracy of 83.91%. URLNet achieved an overall accuracy of 97.07%, falling short of our model’s 98.57%.. These results demonstrate the robustness and effectiveness of our method in complex multi-class classification tasks, indicating its potential as a promising solution for malicious URL detection in cybersecurity.
@@ V-C Cross-dataset testing
To assess the model’s performance and generalization across different phishing datasets, and to identify potential weaknesses or biases within the model, we conducted a cross-dataset testing experiment. Specifically, we used the best-performing pre-trained model fine-tuned on the Grambeddings dataset and directly tested it on the Kaggle binary classification dataset. The results, as shown in Table VI, indicating that URLNet performs well, but Grambeddings has lower accuracy and AUC, with a high false positive rate (FPR) of 0.8667 despite having better recall.
In contrast, our model shows outstanding performance across all metrics, including high precision, recall, an F1 score close to 0.91, and a high AUC. Moreover, it excels in FPR, at 0.0185, surpassing both URLNet and Grambeddings. This is crucial in phishing URL detection models, as it emphasizes the need for low FPR to reduce false alarms in security services, ensuring benign URLs are not misclassified as phishing sites.

### 06 Dec 2023 v2

# PyraTrans: Attention-Enriched Pyramid Transformer for Malicious URL Detection

@@ Abstract
Although advancements in machine learning have driven the development of malicious URL detection technology, current techniques still face significant challenges in their capacity to generalize and their resilience against evolving threats. In this paper, we propose PyraTrans, a novel method that integrates pretrained Transformers with pyramid feature learning to detect malicious URL. PyraTrans utilizes a pretrained CharBERT as its foundation and is augmented with three interconnected feature modules: 1) Encoder Feature Extraction, extracting multi-order feature matrices from each CharBERT encoder layer; 2) Multi-Scale Feature Learning, capturing local contextual insights at various scales and aggregating information across encoder layers; and 3) Spatial Pyramid Attention, focusing on regional-level attention to emphasize areas rich in expressive information. The proposed approach addresses the limitations of the Transformer in local feature learning and regional relational awareness, which are vital for capturing URL-specific word patterns, character combinations, or structural anomalies. In several challenging experimental scenarios, the proposed method has shown significant improvements in accuracy, generalization, and robustness in malicious URL detection. For instance, it achieved a peak F1-score improvement of 40% in class-imbalanced scenarios, and exceeded the best baseline result by 14.13% in accuracy in adversarial attack scenarios. Additionally, we conduct a case study where our method accurately identifies all 30 active malicious web pages, whereas two pior SOTA methods miss 4 and 7 malicious web pages respectively. Codes and data are available at: https://github.com/Alixyvtte/PyraTrans.
@@
Cybersecurity, Phishing Detection, Transformer, Multi-Scale Learning, Pyramid Attention
@@ I Introduction
Recently, pretrained Transformer frameworks, such as BERT [15], have demonstrated promising prospects in various fields including natural language understanding, computer vision, and bioinformatics. This novel computational architecture and training paradigm offer new opportunities in malicious URL detection [16, 17]. They not only have stronger contextual learning capabilities but are also purely data-driven, end-to-end models. Table I presents an example of a BERT token sequence generated from a URL. However, despite the success of pretrained Transformer across multiple disciplines, their technological advantages have yet to be fully explored and manifested in malicious URL detection. Additionally, when applying Transformers to URLs, extra design considerations are essential for local feature learning and focused attention on key segments, as these aspects are crucial for capturing specific word patterns, character combinations, or structural anomalies in URLs.
@@
In this study, we propose PyraTrans, a novel appraoch developed on the foundation of pretrained Transformers, aimed at effectively detecting malicious URLs. This method exploits the various encoding layers of pre-trained Transformers and amalgamates them with multi-scale feature extraction and pyramid attention. Our approach, when tested across a broad range of public benchmarks datasets, exhibited superior performance over previous methods, proving to be more generalizable and robust, particularly in challenging situations like adversarial attacks, cross-dataset evaluation, class-imbalanced learning. We also conduct a case study to evaluate the model in real-world applications.
@@
  • Our method establishes new SOTA performance in malicious URL detection. Throughout extensive experimental evaluations, it consistently demonstrates superior generalization and robustness compared to previous methods. Additionally, it excels in learning under class imbalance and shows improved multi-classification performance, all while reducing the reliance on extensive training data.
  • This work introduces a straightforward multi-order URL feature generation approach by leveraging the multi-level encoding structure of Transformers. Additionally, we develop a multi-scale feature learning module to address the limitations of Transformers in extracting local string pattern from URLs.
  • We combine Transformers with spatial pyramid attention in our method, enabling it to assign regional-level attention. This allows our model to focus on segments with high information density in URLs while reducing the impact of noisy parts, such as random strings.
  • Our research reveals that even previously most popular methods, when subjected to comprehensive evaluations and practical case studies, can exhibit significant limitations. This underscores the critical need for robust case studies in assessing real-world applicability and for a more nuanced evaluation system that considers a method’s generalization and robustness.
@@ IV Methodology
The proposed method employs the CharBERT network structure as the backbone network, integrating an encoder feature extraction module, a multi-scale learning module, and a Spatial Pyramid Attention module.
@@
  • Backbone Network: The CharBERT model improves URL data interpretation and analysis with its advanced subword and character-level embedding which extends based on BERT.
  • Encoder Feature Extraction: This module extracts multi-layer encoder features from CharBERT, aiming to capture URL representations ranging from low-level to high-level.
  • Multi-scale Learning: The module conducts multi-scale local information extraction from multi-layer encoder features and captures the relational information between different encoder feature layers.
  • Spatial Pyramid Attention: This module differentially weights different regions of the feature, highlighting local spatial correlations, allowing flexible focus on information-rich segments in URLs. This contrasts with Transformer’s Multi-Head Attention, which prioritizes positional relationships.
@@
Figure 3: PyraTrans: Composed of Four Core Components. CharBERT, the backbone network for learning character and subword-level features; Encoder Feature Extractor, Multi-Scale Feature Learning, and Spatial Pyramid Attention module for acquiring multi-order, multi-scale, attention-weighted features.
@@ IV-A Backbone Network
We utilize CharBERT[28], an extension of BERT integrating the Transformer architecture with a dual-channel framework, as our pretrained backbone network to capture both subword and character-level features. CharBERT capture character-level information in token sequences through two modules: (1) the Character Embedding Module, encoding character sequences from input tokens and (2) the Heterogeneous Interaction Module, combines features from both character and subword channels, and then independently separates them into distinct representations as input for the encoder layer.
The character-aware embedding of each token is primarily generated through two components: the encoding of individual characters and subword units. These two components are integrated via a dual-channel architecture. To establish contextual character embeddings, we utilize a bidirectional Gated Recurrent Unit (BiGRU) layer. The BiGRU employs a bidirectional recurrent neural network with only the input and forget gates[29]. The architecture diagram of the BiGRU is depicted in Figure 2.
@@ IV-B Encoder Feature Extraction
Although each layer in the BERT family of models takes the output features of the previous layer as input for computation, multiple intricate calculations within each layer’s processing may still result in potential degradation of lower-level and mid-level features, which is detrimental to the complete feature learning process. Li et al. [32] use feature concatenation to integrate aspect features from each layer of BERT for aspect term sentiment classification. This approach, instead of relying only on the final layer for classification features, effectively enhances classification performance by leveraging the distinct features learned at each layer of BERT.
@@
Similar to prior research, we extract outputs from each encoding layer in CharBERT. However, instead of concatenating these layer-wise features, we reorganize them into a higher-dimensional matrix. In this restructured feature matrix, layers function akin to channels in an image. The feature process is as follows: Consider a sequence of outputs $k_{1},k_{2},...,k_{n}$ and $u_{1},u_{2},...,u_{m}$, where each output $k_{i}$ and $u_{j}$ has a rank of $(H,W,C)$, representing the outputs of CharBERT’s word-level and character-level encoders at various layers. $H$ is the batch size, $W$ is the fixed URL sequence length (200 in our model), and $C$ is a 768-dimensional vector for each merged hidden layer output in CharBERT. For example, $k_{1}$ and $u_{2}$ be two tensors representing the sequence and character embeddings, respectively. Let $w$ be the sequence length, and $d$ be the embedding dimension:
@@ IV-C Multi-scale Learning
In extracting URL embedding features, the standard architecture of the Transformer model lacks specialized design for capturing local features. This means it can comprehend the context of the entire input sequence but may not focus on local details. However, capturing local features in URLs is crucial, especially in applications like security analysis, fraud detection, or content categorization. Local features, including specific word patterns, character combinations, or structural anomalies, can be indicative of the nature and intent of a URL.
To more effectively capture these local features, we augment our architecture with multi-scale feature learning, specifically designed for local feature extraction. This module is based on depthwise separable convolutions (DSConv) [33], offering reduced floating-point operations and enhanced computational efficiency. We employ dilated convolutions with varying dilation rates to capture multi-scale information, serving as fundamental operators for expanding the network’s depth and breadth. Formally, the high-dimensional feature represented by outputs from multiple encoder layers is denoted as $M\in\mathbb{R}^{C\times H\times W}$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width. The process begins by applying a single DSConv (conv3 x 3) to $M$, extracting common information denoted as $F_{0}$ for each branch. Specifically:
@@
IV-D Spatial Pyramid Attention
In the multi-scale learning module, a simple element-wise summation of features from different scales may inadvertently downplay the importance of informative branches, treating all scales equally. Additionally, while Transformers offer token-level attention, for URL feature learning, regional-level attention is crucial due to the presence of distinct information-dense areas (like domain names) and regions with noise (such as random parameters) in URLs. To address these issue, we integrate a Spatial Pyramid Attention module following our multi-scale learning [35].
@@ V Experiments
    Assess stacked feature layer configurations for multi-layer extraction.
@@
    Explore data scale dependency with varied training dataset sizes.
@@
    Evaluate model generalization across datasets.
@@
    Multi-class classification.
@@
    Test robustness using adversarial samples.
@@
    Evaluate practicality with recent active malicious URLs.
Setup. The pre-trained CharBERT was trained on the English Wikipedia corpus, consisting of a total of 12GB and approximately 2,500 million words. Hyperparameter tuning during fine-tuning led to a batch size of 64, AdamW optimizer (initial learning rate: 2e-5, weight decay: 1e-4), 0.1 dropout rate, and 5 training epochs. We utilized PyTorch 2.0, NVIDIA CUDA 11.8, and Python 3.8, conducting training on NVIDIA A100 GPUs. The final model for each experiment was selected based on the best validation loss.
@@
V-A Evaluate the Effectiveness of Multi-level Features
We develop a multi-layer feature extraction module to distill semantic features from different layers of the Transformer encoders in our backbone network. Our approach involves stacking embedding outputs from different layers to form a multi-layered feature matrix, which we optimize to enhance URL semantic feature representation. Through this, we investigate the complexity of feature representation across the network and aim to demonstrate the effectiveness of multi-layered features in improving overall model performance.
We create a training corpus from the GramBeddings dataset, randomly selecting 128k URLs from the 80k available in the training set while maintaining the proportion of malicious and benign URLs. We also randomly sample 32k URLs for testing and validation respectively. The performance of various stacked configurations and the incremental gains are shown in Table IV. We observe a consistent improvement in evaluation metrics as we incrementally incorporate additional layers, namely the last 2, 3, 4, and 5 layers. Ultimately, the stacking of 12 embedding output layers achieves the best performance in URL detection tasks, demonstrating the effectiveness of integrating both lower and deeper layers in the model architecture.
V-B Comparison with Baselines
In this section, we compare the performance of PyraTrans with baselines using binary and multi-class malicious URL detection.
@@ V-B1 Binary classification
As shown in Table V, our method exhibits significant advantages in class-imbalanced scenarios compared to other approaches. With just 1% of training data, our model achieves an accuracy of 0.9837, with a Precision of 0.7632, surpassing the best baseline Precision of 0.1808. The F1 score of our method is four times higher than the best baseline performance. As training data increases, our F1 score reaches 0.7027, substantially exceeding the baseline peak of 0.5515, with a Recall 15.3 higher than the best baseline result. These experiments demonstrate the substantial improvements our method brings in identifying malicious URL samples, significantly reducing false positives. Moreover, the high AUC reflects our model’s confidence in the identified samples, indicating its accurate capture of key differences between malicious and benign samples.
We observe that while URLNet achieves high accuracy on larger datasets, it suffers from a high false negative rate, particularly on smaller datasets. Notably, URLNet fails to identify any malicious URLs when the training data size is reduced to below 5%. In contrast, Grambeddings shows significant sensitivity to data, with its performance in detecting malicious URLs varying greatly across different training sample sizes. For instance, at 60% training data, its Precision drops to 0.6137, and at 40%, its Recall falls to just 0.0817. Compared to these, our method demonstrates consistent and reliable performance across various data scales, indicating its robustness to small-scale and class-imbalanced data scenarios.
V-B2 Multi-classification
@@
Fig.6 illustrates the performance of our method and the baseline methods across all four categories. Our method surpasses the baseline methods in each metric. The average ROC curve highlights our method’s efficacy with a TPR of almost 90% at a low FPR of 0.001, surpassing other methods that achieve around 75%. Grambeddings struggled in recognizing negative samples from various categories, resulting in a 50% F1 score for defacement and phishing URLs and an overall accuracy of 83.91%. URLNet achieved an overall accuracy of 97.07%, falling short of our model’s 98.57%. These results demonstrate the robustness and effectiveness of our method in complex multi-class classification tasks, indicating its potential as a promising solution for malicious URL detection in cybersecurity.
@@ V-C Cross-dataset Testing
To evaluate the generalization of models and to identify any potential weaknesses or biases, we set up a cross-dataset testing experiment. We do this by training the model on the Grambeddings dataset and subsequently testing it on the Kaggle binary classification dataset. It is noteworthy that the Grambeddings and Kaggle datasets significantly differ in their data collection times and sources.
The results, as shown in Table VI, indicate a marked decline in performance of baseline methods on data not included in their training set, with URLNet’s accuracy dropping to 0.8823 and Grambeddings’ accuracy reducing to 0.5214 and 0.4647, respectively. In contrast, our model maintained high accuracy on external datasets, achieving an AUC of 0.9705. This underscores that the knowledge gained from one dataset by our method can be effectively generalized to others, even if the data was collected much later than our model’s pretraining period. This demonstrates the adaptability and long-term applicability of our approach to different data environments.